# Probabilistic Consistency in Machine Learning and Its Connection to Uncertainty Quantification

### v1

@@ Abstract
Machine learning (ML) is often viewed as a powerful data analysis tool that is easy to learn because of its black-box nature. Yet this very nature also makes it difficult to quantify confidence in predictions extracted from ML models, and more fundamentally, to understand how such models are mathematical abstractions of training data. The goal of this paper is to unravel these issues and their connections to uncertainty quantification (UQ) by pursuing a line of reasoning motivated by diagnostics. In such settings, prevalence – i.e. the fraction of elements in class – is often of inherent interest. Here we analyze the many interpretations of prevalence to derive a level-set theory of classification, which shows that certain types of self-consistent ML models are equivalent to class-conditional probability distributions. We begin by studying the properties of binary Bayes optimal classifiers, recognizing that their boundary sets can be reinterpreted as level-sets of pairwise density ratios. By parameterizing Bayes classifiers in terms of the prevalence, we then show that they satisfy important monotonicity and class-switching properties that can be used to deduce the density ratios without direct access to the boundary sets. Moreover, this information is sufficient for tasks such as constructing the multiclass Bayes-optimal classifier and estimating inherent uncertainty in the class assignments. In the multiclass case, we use these results to deduce normalization and self-consistency conditions, the latter being equivalent to the law of total probability for classifiers. We also show that these are necessary conditions for arbitrary ML models to have valid probabilistic interpretations. Throughout we demonstrate how this analysis informs the broader task of UQ for ML via an uncertainty propagation framework.
@@ 1 Introduction
From our perspective, the formulation of a complete uncertainty quantification (UQ) framework for machine learning (ML) remains an unresolved problem. Several works have addressed empirical aspects of this problem through the lens of Monte Carlo methods and related sampling techniques [1, 2, 3, 4]. Moreover, it is well known that classification theory in general can be grounded in probability [5]. Indeed, this perspective has led to important observations, such as the recognition that the Bayes classifier, which minimizes the expected classification error, also arises from optimization of surrogate loss functions [6, 7, 8, 9]. In this context, many authors have also addressed elements of uncertainty estimation as a task in calibration, i.e. ensuring that probabilistic ML outputs are consistent with distributional information about data [10, 11, 12, 13, 14, 15, 16]. However, these works have not considered more fundamental and perhaps subtle questions: (i) are ML predictions always mathematically consistent with probability theory; (ii) can one meaningfully associate aleatoric uncertainty estimates with arbitrary ML models [17]; and thus, (iii) do they even permit meaningful UQ? Stated differently, when is calibration possible, and what are the effects that would inform the calibration process?
@@
A key difficulty in answering these questions arises from the fact that it is not always clear how arbitrary ML models are abstract representations of training data. For example, the Bayes classifier is known to be a manifestation of underlying conditional distributions $\Pr[r|C]$ of “inputs” $r$ generated by class $C$. But by the same token, it is not obvious that such classifiers contain all of the information about these distributions. Thus, it is possible that certain ML models can provide minimum error guarantees in a global sense,***That is, by minimizing the expected classification error. but not pointwise estimates of their confidence. More generally, if an arbitrary classifier is to admit meaningful UQ, then we assert that it must be equivalent in some sense to the underlying conditional distributions $\Pr[r|C]$.
@@
The purpose of this manuscript is to unravel such issues by deriving a level-set theory of classification that establishes necessary conditions for there to be a mathematical correspondence between classifiers $\hat{C}$ and the distributions $\Pr[r|C]$. Importantly, this correspondence permits a deeper analysis of UQ for ML in terms of a formal uncertainty propagation framework and thereby facilitates tasks such as estimation of inherent uncertainty (a variation on aleatoric uncertainty) in the class labels. The main idea of this analysis is to study the properties of binary Bayes classifiers, and in particular their boundary sets, which we reinterpret as level sets of pairwise density ratios quantifying the likelihood that an input $r$ arises from a given class. By parameterizing these level sets in terms of the prevalence (fraction of elements in a class), we arrive at an important montonicity property that holds in general for Bayes classifiers. Without having explicit access to the boundary sets, we then use this property to deduce the density ratios, and hence information about $\Pr[r|C]$, in terms of the prevalence value at which the class changes. In extending this result to the multiclass setting, we also identify normalization and self-consistency criteria, the latter expressing the law of total probability for classifiers. By recourse to the density ratios, our analysis then: (i) demonstrates the sense in which monotonicity, self-consistency, and normalizability are necessary conditions for establishing the correspondence between $\hat{C}$ and $\Pr[r|C]$; and (ii) shows how the aforementioned analyses can account for distribution shifts due to changes in prevalence.
@@
First and most simply, Eq. (1) allows one to better identify and study individual sources of uncertainty. Note, for example, that even when $\Pr[r|C]$ and $\Pr[C]$ are known exactly, it is possible that $0<\Pr[C|r]<1$, in which case $r$ cannot unambiguously be assigned a single class. We refer to this effect as “inherent” uncertainty because it is a fundamental property of the underlying input space [2]. In contrast, the act of choosing a hypothesis set ${\mathcal{H}}$ of classifiers is equivalent to selecting a family of distributions $\Pr[r|C]$ for regresssion. This choice may not be well adapted to the data and thereby introduce additional uncertainty into $\Pr[r|C]$, in this case associated with modeling per se. In classification theory, these issues are sometimes acknowledged in the context of $H$-consistency bounds [6, 7, 8, 18], but they have not been fully separated from estimates of $\Pr[C]$, which is yet a third and often significant source of uncertainty. Thus, Eq. (1) allows us to disentangle such effects via the formal uncertainty propagation framework, which quantifies their individual and compounding contributions to the total uncertainty budget.
@@
Finally, a historical note is in order. A variation on the monotonicity property was previously noticed as far back as 2005 by Langford and Zadrozny [20], primarily in the context of binary classifiers. To a lesser extent Ref. [32] examined similar ideas, and Ref. [33] considered some extensions to the multiclass setting. However, our analysis more thoroughly examines the properties of the Bayes classifiers and thereby discovers that the concept of monotone classifier requires significant refinement and generalization. This is necessary to deduce the self-consistency and normalizability criteria, which appear to have been acknowledged but not fully understood in previous works; see, e.g. the Discussion section in Ref. [32]. In a related vein, extensions to the multiclass setting require extreme care due to issues associated with non-overlapping supports of the conditional distributions. In fact, this leads to a fundamental issue of how to interpret classifiers that are evaluated on points outside of their domain of action, and we show that this is analogous to trying to evaluate indeterminate forms such as $/0$ 0. Thus, our analysis is partly meant to address unresolved mathematical questions in previous work.
The rest of this manuscript is organized as follows. Section 2 gives a broad overview of our perspective on the relationship between classification and probability (2.1), which allows us to ground the present manuscript in the broader UQ literature and identify sources of uncertainty in ML models (2.2). Section 3 provides background theory for the binary setting. Section 4 presents our main results, which fully generalize the level-set and class-switching theorems to an arbitrary multi-class settings. Section 5 provides examples of our main results and uses these to highlight practical issues of determining when a classifier is self-consistent. Section 6 considers our results more broadly in the context of ML theory and discusses limitations and open directions.
@@ 1.1 On the Conventions Used Herein
[…]
@@ 2.1 Overview of our perspective: what is a classifier?
To elaborate on this point, consider that Eq. (1) quantifies all of the objective information available for making a decision, given a fixed space $\Gamma$ and the probabilities $\Pr[r|C]$ and $\Pr[C]$. How one chooses to use that information depends on the application at hand; e.g. a Bayes optimal classifier may not be suitable for settings that heavily penalize one type of misclassification over another. But in light of Eq. (1), one could argue that $\Pr[r|C]$ and $\Pr[C]$ suffice for constructing “reasonable” classifiers and estimating their induced uncertainties. In this vein, our analysis ultimately seeks to answer the question: given a hypothesis set ${\mathcal{H}}$ of classifiers that are “adapted” to the data, can we deduce $\Pr[r|C]$? Were this possible, the act of training a classifier would be equivalent to statistical regression, in which case it would be possible to not only estimate uncertainties, but also further modify the classifier to account for distributions shifts associated with changes in $\Pr[C]$. Thus, the main task of this manuscript can be restated as determining the sense in which the function $\hat{C}$ models the probability $\Pr[r|C]$.
To understand why this is so difficult, it is instructive to examine the ways in which classifiers in the spirit of Eq. (3) fail to represent probabilities. Under the best of circumstances, we would want to construct $\hat{C}$ such that
@@
Let ${{\rm X}^{K}}$ be the probability simplex, and let $\hat{C}:\Gamma\times{{\rm X}^{K}}\to{\mathbb{K}}$. For $k<k^{\prime}$ and $\alpha\in{{\rm X}^{2}}$, let $q(\alpha;k,k^{\prime}):{{\rm X}^{2}}\to{{\rm X}^{K}}$ be the vector whose elements are $q(\alpha;k,k^{\prime})_{k}=\alpha_{1}$, $q(\alpha;k,k^{\prime})_{k^{\prime}}=\alpha_{2}=1-\alpha_{1}$, and $q(\alpha;k,k^{\prime})_{j}=0$ for $j\neq k,k^{\prime}$. Then we say that $\hat{C}$ is a monotone classifier if for every pair $k,k^{\prime}$, $k<k^{\prime}$:
@@
While definition 1 seems abstract, its interpretation is straightforward in the context of a binary classifier. In this case, $\Gamma_{k,k^{\prime}}=\Gamma$, and $\hat{C}$ reduces to a function of the form $\hat{C}(r,(\alpha_{1},1-\alpha_{1}))$ for $\alpha_{1}\in[0,1]$. Figure 1 shows a simple monotone classifier sampled on a discrete grid of $\alpha_{1}$ values. In fact, these are Bayes-optimal classifiers associated with populations $\Omega(\chi)$ having test prevalence values $\chi=q=(\alpha_{1},1-\alpha_{1})$ [see Sec. 5 for details]. This example illustrates an intuitive and powerful property: monotone classifiers have a set inclusion (or monotone class [43]) structure parameterized by $\alpha$, and as a result, the assigned class of each point changes at one and only one value of $\alpha$. Key tasks in the next sections are to (i) demonstrate that Bayes classifiers satisfy Def. 1, (ii) prove that the value of $\alpha$ at which the class changes determines $\Pr[r|C]$; and (iii) show that the sets $\Gamma_{k,k^{\prime}}$ are tied to the supports of $\Pr[r|C]$. But to arrive at this point, we must formalize the monotone structure in terms of set theory; see Sec. 4. For now it is sufficient to note that in the context of Fig. 1, we may view $q$, as a parameter that “deforms one classifier into another.” It is for this reason that we refer to this variable as an affine prevalence, i.e. to highlight its connection to geometry.§§§We believe that the connection between our work and differential geometry runs much deeper than initially meets the eye. The terminology “affine prevalence” is also meant to suggest the idea that the relative-probability level-sets discussed in the next sections may have a viable interpretation as tangent spaces connected by $q$. See Sec. 6, as well as the curves on Fig. 1. Note that $q\in{{\rm X}^{K}}$ should be treated as a variable, not a fixed parameter.
@@
Figure 1: Example of a binary monotone classifier. The color scale denotes the probability densities of two Gaussian distributions quantifying the probabilities $\Pr[r|C]$ for two classes. The top (red) density is associated with class 111, and the bottom (blue) density is associated with class 222. The hyperbolas are classification boundaries for different values of the affine prevalence $q=(\alpha_{1},1-\alpha_{1})$. Each boundary corresponds to a different classifier. For the plotted values of $\alpha_{1}\geq 0.05$, the points on the same side as the foci are assigned to class 222, and those on the other side are assigned to class 111. For $\alpha_{1}=10^{-3}$, the assignment is reversed. These boundaries are also Bayes-optimal for test populations $\Omega(\chi)$ having test prevalence values $\chi=q$. [Shading does not correspond to any particular test prevalence.] Note that as a function of $\alpha_{1}$, the point $x$ changes class one and only one time. In fact, $\hat{C}(r,q)$ is a monotone decreasing function of $\alpha_{1}$. The classification boundaries also suggest an important set inclusion structure associated with the classification domains. See the top-right panel of Fig. 3 (which shows the same distributions) and Eqs. (41)–(42c) for more details.
@@ 2.2 Implications for UQ
Momentarily assuming that the claims of Sec. 2.1 are true, we pause to consider how our perspective informs the broader task of UQ for classification and ML. The purposes of this discussion are twofold. First, our analysis largely considers classification in the idealized limit of infinite data and perfect optimization. Distinguishing sources of uncertainty helps to clarify the assumptions used herein. Second, this idealized limit yields the smallest-possible uncertainty in the class labels, upon which all other effects build. We refer the reader to Refs. [1, 19, 2, 3] and the references therein for other perspectives on UQ for ML.
@@
$\tilde{u}(r)=1-\max_{C}\left\{\frac{\left[\Pr[C]+\epsilon_{\chi}\right]\left[\Pr[r|C]+\epsilon_{\rm dist}\right]}{\sum_{C^{\prime}}\left[\Pr[C^{\prime}]+\epsilon_{\chi}\right]\left[\Pr[r|C^{\prime}]+\epsilon_{\rm dist}\right]}\right\}+\epsilon_{\rm num},$ (8)
where $\epsilon_{\chi}$ quantifies uncertainty in the prevalence estimate, $\epsilon_{\rm dist}$ quantifies uncertainty associated with the conditional probability distributions, and $\epsilon_{\rm num}$ quantifies other sources of uncertainty that may be associated with numerical realizations of $u(r)$.
To understand how sources of uncertainty contribute to $\epsilon_{\chi}$ and $\epsilon_{\rm dist}$, consider the typical steps in a classification process. Often one begins by collecting training and test data. Several authors have demonstrated that the test prevalence $\chi=\Pr[C]$ can be objectively estimated without recourse to classification, which is the so-called quantitation or quantification problem [22, 23, 24, 25, 26, 27, 28, 29, 30]. Moreover, it is well known that the amount of training and test data controls the uncertainty $\epsilon_{\chi}$ in such estimates, so that we refer to this effect as sampling uncertainty. The next step is often to choose a hypothesis set ${\mathcal{H}}$ of functions that in principle include the classifier of interest. By this, we mean the hypothesis set is assumed to contain a $\hat{C}^{\star}$ that minimizes a chosen loss function $\mathcal{L}$. In this work, we consider the 0111 loss, so that $\hat{C}^{\star}$ is the Bayes optimal classifier. However, given finite training data, one must approximate $\mathcal{L}$ in terms of an empirical loss function $\hat{\mathcal{L}}$ whose minimizer may differ from $\hat{C}^{\star}$, even when $\hat{C}^{\star}\in{\mathcal{H}}$. This is another manifestation of sampling uncertainty, and this is equivalent to errors induced in fitting distributions for $\Pr[r|C]$ to finite data. Such effects contribute to $\epsilon_{\rm dist}$.
@@
Figure 2 shows how these sources of uncertainty are connected to one another and propagate through a fishbone diagram, which yields the total uncertainty in the class labels. Note in particular that construction of the conditional distributions $\Pr[C|r]$ is equivalent to estimating the inherent uncertainty, assuming nothing else contributes. This figure, as well as Eq. (8), also clarifies our key perspective. If we cannot unambiguously associate a classifier $\hat{C}$ with the conditional distributions $\Pr[r|C]$, it is not possible to even compute $\tilde{u}(r)$. Moreover, this association cannot be made in an ad hoc or post-hoc manner. The process of training a classifier itself can induce probabilistic information in $\hat{C}$, and auxiliary modeling choices on top of this can lead to unquantifiable model-form errors or worse, inconsistency with probability. We also observe that because $\epsilon_{\chi}$ depends primarily on sampling, it is in generally a source of epistemic (i.e. reducible) uncertainty. Thus, $\Pr[r|C]$ fully determines the aleatoric or irreducible uncertainty, and our main task amounts to determining when this concept can be meaningfully applied to a classifier.

### v2

@@ Abstract
Machine learning (ML) is often viewed as a powerful data analysis tool that is easy to learn because of its black-box nature. Yet this very nature also makes it difficult to quantify confidence in predictions extracted from ML models, and more fundamentally, to understand when they admit probabilistic interpretations. The goal of this paper is to unravel these issues and their connections to uncertainty quantification (UQ) by deriving a level-set theory of classification that establishes an equivalence between certain types of self-consistent ML models and class-conditional probability distributions. We begin by studying the properties of binary Bayes classifiers, recognizing that their boundary sets can be reinterpreted as level-sets of density ratios quantifying the relative probability that a sample point belongs to a given class. By promoting the prevalence (fraction of elements in a class) to the role of an affine parameter that orders these level sets, we show that Bayes classifiers satisfy important monotonicity and class-switching properties that can be used to extract the density ratios without direct access to the boundary sets. Moreover, this information is sufficient for tasks such as constructing the multiclass Bayes classifier from its pairwise counterparts and estimating inherent uncertainty in the class assignments. In the multiclass case, we use these results to derive normalization and self-consistency properties of Bayes classifiers, the latter being equivalent to the law of total probability. We then show how these properties equip arbitrary ML models with valid probabilistic interpretations arising from inherent class conditional probability distributions. Throughout, we demonstrate how this analysis informs the broader task of UQ for ML via an uncertainty propagation framework.
@@ 1 Introduction
From our perspective, the formulation of a complete uncertainty quantification (UQ) framework for machine learning (ML) remains an unresolved problem. Several works have addressed empirical aspects of this problem through the lens of Monte Carlo methods and related sampling techniques [1, 2, 3, 4]. Moreover, it is well known that classification theory can be grounded in probability [5]. Indeed, this perspective has led to important observations, such as the recognition that the Bayes classifier, which minimizes the expected classification error, also arises from optimization of surrogate loss functions [6, 7, 8, 9]. In this context, many authors have addressed elements of uncertainty estimation as a task in calibration, i.e. ensuring that probabilistic ML outputs are consistent with distributional information about data [10, 11, 12, 13, 14, 15, 16]. However, these works have not considered more fundamental and perhaps subtle questions: (i) are ML predictions always mathematically consistent with probability theory; (ii) can one meaningfully associate aleatoric uncertainty with arbitrary ML models [17]; and thus, (iii) do they even permit meaningful UQ? Stated differently, when is calibration possible, and what mathematical properties of classifiers should inform calibration?
@@
A key difficulty in answering these questions arises from the fact that it is not always clear how, or even if, arbitrary ML models are abstract representations of training data. For example, a discriminative Bayes classifier by definition expresses information about the underlying conditional probabilities $\Pr[r|C]$ of “inputs” $r$ generated by elements from class $C$. But while such classifiers provide minimum error guarantees in a global sense,***That is, by minimizing the expected classification error. they generally do not yield a priori pointwise estimates of their confidence. This begs the question of whether discriminative classifiers contain all of the information about the distributions $\Pr[r|C]$, or if something has been lost. Such observations also lead us to speculate that if an arbitrary classifier is to admit meaningful UQ, then it must be “fully equivalent” to the underlying conditional distributions $\Pr[r|C]$.
@@
The purpose of our manuscript is to understand in what sense this equivalence between classifiers and distributions holds. The main idea of our analysis is to first determine which properties of Bayes classifiers arise from their probabilistic structure and then demonstrate how these properties alone enable one to associate an arbitrary classifier $\hat{C}$ with distributions $\Pr[r|C]$. To achieve this in practice, we consider the boundary sets of Bayes classifiers, which are reinterpreted as level sets of pairwise density ratios quantifying the likelihood that an input $r$ arises from a given class. By promoting prevalence (fraction of elements in a class) to the role of an affine parameter that orders these level sets, we arrive at an important montonicity property that holds in general for Bayes classifiers. We then demonstrate how this property can be used to deduce the density ratios, and hence information about $\Pr[r|C]$, in terms of the affine prevalence value at which the class changes. In extending this result to the multiclass setting, we also identify normalization and self-consistency criteria, the latter expressing the law of total probability for classifiers. By recourse to the density ratios, our analysis then: (i) demonstrates the sense in which monotonicity, self-consistency, and normalizability enable one to establish a correspondence between arbitrary $\hat{C}$ and $\Pr[r|C]$; and (ii) shows how the aforementioned analyses can account for distribution shifts due to changes in the prevalence of a test population.
@@
First and most simply, Eq. (1) allows one to better identify and study individual sources of uncertainty. Note, for example, that even when $\Pr[r|C]$ and $\Pr[C]$ are known exactly, it is possible that $0<\Pr[C|r]<1$, in which case $r$ cannot unambiguously be assigned a single class. We refer to this effect as “inherent” uncertainty because it is a fundamental property of the underlying input space [2]. In contrast, the act of choosing a hypothesis set $\mathcal{H}$} of classifiers is equivalent to selecting a family of distributions $\Pr[r|C]$ for regresssion. This choice may not be well adapted to the data and thereby introduce additional uncertainty into $\Pr[r|C]$, in this case associated with modeling per se. In classification theory, these issues are sometimes acknowledged in the context of $H$-consistency bounds [6, 7, 8, 18], but they have not been fully separated from estimates of $\Pr[C]$, which is yet a third and often significant source of uncertainty. Thus, Eq. (1) allows us to disentangle such effects via a formal uncertainty propagation framework, which quantifies their individual and compounding contributions to the total uncertainty budget.
@@
Finally, a historical note is in order. A variation on the monotonicity property was previously noticed as far back as 2005 by Langford and Zadrozny [20], primarily in the context of binary classifiers. To a lesser extent Ref. [32] examined similar ideas, and Ref. [33] considered some extensions to the multiclass setting. However, our analysis more thoroughly examines the properties of the Bayes classifiers and thereby discovers that the concept of monotone classifier requires significant refinement and generalization. This is necessary to deduce the self-consistency and normalizability criteria, which appear to have been acknowledged but not fully understood in previous works; see, e.g. the Discussion section in Ref. [32]. In a related vein, extensions to the multiclass setting require extreme care due to issues associated with non-overlapping supports of the conditional distributions. In fact, this leads to a fundamental issue of how to interpret classifiers that are evaluated on points outside of their domain of action, and we show that this is analogous to trying to evaluate indeterminate forms such as $0/0$. Thus, our analysis is partly meant to address unresolved mathematical questions in previous work.
The rest of this manuscript is organized as follows. Section 2 gives a broad overview of our perspective on the relationship between classification and probability (2.1), which allows us to ground the present manuscript in the broader UQ literature and identify sources of uncertainty in ML models (2.2). Section 3 provides background theory for the binary setting. Section 4 presents our main results, which fully generalize the level-set and class-switching theorems to an arbitrary multi-class settings. Section 5 provides examples of our main results and uses these to highlight practical issues of determining when a classifier has a probabalistic interpretation. Section 6 considers our results more broadly in the context of ML theory and discusses limitations and open directions.
@@ 1.1 On the Conventions Used Herein
  • We use the extended number system in which $\infty$ is a valid number (not just a limit) satisfying the properties that: $a\cdot\infty=\infty$ for $a$ positive and bounded; $a/0=\infty$ for $a$ positive and bounded; the product $\infty\cdot 0$ (when interpreted as $\infty\cdot 0^{+}$) is a non-negative indeterminate number; and the ratio $0/0$ (when interpreted as $0^{+}/0^{+}$) is a non-negative, indeterminate number. We also use the notation $0,\infty]$ to denote the interval containing its endpoints, including $\infty$. See also the preface of Ref. [34], for example.
@@ 2.1 Overview of our perspective: what is a classifier?
To elaborate on this point, consider that Eq. (1) quantifies all of the information available for making a decision, given a fixed space $\Gamma$ and the probabilities $\Pr[r|C]$ and $\Pr[C]$. How one chooses to use that information depends on the application at hand; e.g. a Bayes optimal classifier may not be suitable for settings that heavily penalize one type of misclassification over another. But in light of Eq. (1), one could argue that $\Pr[r|C]$ and $\Pr[C]$ suffice for constructing “reasonable” classifiers and estimating their induced uncertainties. In this context, our analysis ultimately seeks to answer the question: given a classifier that is sufficiently “adapted” to the data, can we deduce $\Pr[r|C]$? Were this possible, it would show that the act of training a classifier is equivalent to statistical regression, i.e. a modeling exercise. But more importantly, it would also allow one to modify a classifier to account for distribution shifts (e.g. in the prevalence $\Pr[C]$) and alternative notions of loss. Thus, the main task of our manuscript can be restated as determining the sense in which the function $\hat{C}$ represents the probability $\Pr[r|C]$
To understand why this task is difficult, it is instructive to examine the ways in which classifiers in the spirit of Eq. (3) fail to represent probabilities. Under the best of circumstances, we would want to construct $\hat{C}$ such that
@@
Let ${\rm X}^{K}$} be the probability simplex, and let $\hat{C}:\Gamma\times{{\rm X}^{K}}\to{\mathbb{K}}$. For $k<k^{\prime}$ and $\alpha\in{{\rm X}^{2}}$, let $q(\alpha;k,k^{\prime}):{{\rm X}^{2}}\to{{\rm X}^{K}}$ be the vector whose elements are $q(\alpha;k,k^{\prime})_{k}=\alpha_{1}$, $q(\alpha;k,k^{\prime})_{k^{\prime}}=\alpha_{2}=1-\alpha_{1}$, and $q(\alpha;k,k^{\prime})_{j}=0$ for $j\neq k,k^{\prime}$. Then we say that $\hat{C}$ is a monotone classifier if for every pair $^{\prime}, k<kk<k^{\prime}$:
@@
While definition 1 seems abstract, its interpretation is straightforward in the context of a binary classifier. In this case, $\Gamma_{k,k^{\prime}}=\Gamma$, and $\hat{C}$ reduces to a function of the form $\hat{C}(r,(\alpha_{1},1-\alpha_{1}))$ for $\alpha_{1}\in[0,1]$. Figure 1 shows a simple monotone classifier sampled on a discrete grid of $\alpha_{1}$ values. In fact, these are Bayes-optimal classifiers associated with populations $\Omega(\chi)$ having test prevalence values $\chi=q=(\alpha_{1},1-\alpha_{1})$ [see Sec. 5 for details]. This example illustrates an intuitive and powerful property: monotone classifiers have a set inclusion (or monotone class [43]) structure parameterized by $\alpha$, and as a result, the assigned class of each point changes at one and only one value of $\alpha$. Key tasks in the next sections are to (i) demonstrate that Bayes classifiers satisfy Def. 1, (ii) prove that the value of $\alpha$ at which the class changes determines $\Pr[r|C]$; and (iii) show that the sets $\Gamma_{k,k^{\prime}}$ are tied to the supports of $\Pr[r|C]$. But to arrive at this point, we must formalize the monotone structure in terms of set theory; see Sec. 4. For now it is sufficient to note that in the context of Fig. 1, we may view $q$, as a parameter that “deforms one classifier into another.” It is for this reason that we refer to this variable as an affine prevalence, i.e. to highlight its connection to geometry.§§§We believe that the connection between our work and differential geometry runs much deeper than initially meets the eye. The terminology “affine prevalence” is also meant to suggest the idea that the density ratio level-sets discussed in the next sections may have a viable interpretation as tangent spaces connected by $q$. See Sec. 6, as well as the curves on Fig. 1. Note that $q\in{{\rm X}^{K}}$ should be treated as a variable, not a fixed parameter.
@@
Figure 1: Example of a binary monotone classifier. The color scale denotes the probability densities of two Gaussian distributions quantifying the probabilities $\Pr[r|C]$ for two classes. The top (red) density is associated with class $1$, and the bottom (blue) density is associated with class $2$. The hyperbolas are classification boundaries for different values of the affine prevalence $q=(\alpha_{1},1-\alpha_{1})$. Each boundary corresponds to a different classifier. For the plotted values of $\alpha_{1}\geq 0.05$, the points on the same side as the foci are assigned to class $2$, and those on the other side are assigned to class $1$. For $\alpha_{1}=10^{-3}$, the assignment is reversed. These boundaries are also Bayes-optimal for test populations $\Omega(\chi)$ having test prevalence values $\chi=q$. [Shading does not correspond to any particular test prevalence.] Note that as a function of $\alpha_{1}$, the point $x$ changes class one and only one time. In fact, $\hat{C}(r,q)$ is a monotone decreasing function of $\alpha_{1}$. The classification boundaries also suggest an important set inclusion structure associated with the classification domains. See the top-right panel of Fig. 3 (which shows the same distributions) and Eqs. (41)–(42c) for more details.
@@ 2.2 Implications for UQ
Momentarily assuming that the claims of Sec. 2.1 are true, we pause to consider how our perspective informs the broader task of UQ for classification and ML. The reasons for doing so are twofold. First, we primarily consider classification in the idealized limit of infinite data and perfect optimization. Framing this analysis in the context of UQ helps to distinguish sources of uncertainty and thereby clarify the assumptions used herein. Second, this idealized limit yields the smallest-possible uncertainty in the class labels, upon which all other effects build. Thus, our discussion highlights the fact that the relationship between classifiers and probability distributions is fundamental to UQ of ML. We refer the reader to Refs. [1, 19, 2, 3] and the references therein for other perspectives on this topic.
@@
$\tilde{u}(r)=1-\max_{C}\left\{\frac{\left[\Pr[C]+\epsilon_{\chi}(C)\right]\left[\Pr[r|C]+\epsilon_{\rm dist}(C)\right]}{\sum_{C^{\prime}}\left[\Pr[C^{\prime}]+\epsilon_{\chi}(C)\right]\left[\Pr[r|C^{\prime}]+\epsilon_{\rm dist}(C)\right]}\right\}+\epsilon_{\rm num},$ (8)
where $\epsilon_{\chi}(C)$ quantifies uncertainty in the prevalence estimate as a function of the class $C$, $\epsilon_{\rm dist}(C)$ quantifies uncertainty associated with the conditional probability distributions, and $\epsilon_{\rm num}$ quantifies other sources of uncertainty that may be associated with numerical realizations of $u(r)$.
To understand how sources of uncertainty contribute to $\epsilon_{\chi}$ and $\epsilon_{\rm dist}$, consider the typical steps in a classification process. Often one begins by collecting training and test data. Several authors have demonstrated that the test prevalence $\chi=\Pr[C]$ can be objectively estimated without recourse to classification, which is the so-called quantitation or quantification problem [22, 23, 24, 25, 26, 27, 28, 29, 30]. Moreover, it is well known that the amount of training and test data controls the uncertainty $\epsilon_{\chi}$ in such estimates, so that we refer to this effect as sampling uncertainty. The next step is often to choose a hypothesis set $\mathcal{H}$} of functions that in principle include the classifier of interest. By this, we mean the hypothesis set is assumed to contain a $\hat{C}^{\star}$ that minimizes a chosen loss function $\mathcal{L}$. In this work, we consider the 011 loss, so that $\hat{C}^{\star}$ is the Bayes optimal classifier. However, given finite training data, one must approximate $\mathcal{L}$ in terms of an empirical loss function $\hat{\mathcal{L}}$ whose minimizer may differ from $\hat{C}^{\star}$, even when $\hat{C}^{\star}\in{\mathcal{H}}$. This is another manifestation of sampling uncertainty, and this is equivalent to errors induced in fitting distributions for $\Pr[r|C]$ to finite data. Such effects contribute to $\epsilon_{\rm dist}$.
@@
Figure 2 shows how these sources of uncertainty are connected to one another and propagate through a fishbone diagram, which yields the total uncertainty in the class labels. Note in particular that construction of the conditional distributions $\Pr[C|r]$ is equivalent to estimating the inherent uncertainty, assuming nothing else contributes. This figure, as well as Eq. (8), also clarifies our key perspective. If we cannot unambiguously associate a classifier $\hat{C}$ with the conditional distributions $\Pr[r|C]$, it is not possible to even compute $\tilde{u}(r)$. Moreover, this association cannot be made in an ad hoc or post-hoc manner. The process of training a classifier itself can induce probabilistic information in $\hat{C}$, and auxiliary modeling choices on top of this can lead to unquantifiable model-form errors or worse, inconsistency with probability. We also observe that because $\epsilon_{\chi}$ depends primarily on sampling, it is in generally a source of epistemic (i.e. reducible) uncertainty. Thus, $\Pr[r|C]$ fully determines the aleatoric or irreducible uncertainty, and our main task amounts to determining when this concept can be meaningfully applied to a classifier. For this reason, the remainder of the manuscript primarily focuses on settings in which $u(r)$ is the only source of uncertainty.